OpenStudy (anonymous):

The lengths of the sides of a rectangle are doubled. Which statement is true about the rectangle? Its area and perimeter become double. Its area becomes four times greater, and perimeter becomes double. Its area and perimeter become four times greater. Its area becomes double, and perimeter becomes four times greater.

OpenStudy (pfenn1):

If the area of a rectangle are L xH and the perimeter is 2L + 2H then what if you double the lehgth and height area=2L x 2H perimeter= 2 (2L)+2(2H) so how does the area of the first rectangle compare to the second? And the perimeter?

OpenStudy (anonymous):

does the area double & are 4 x greater?

OpenStudy (pfenn1):

do you mean area doubles and perimeter is 4x greater?

OpenStudy (anonymous):

yes, my badd

OpenStudy (pfenn1):

The area of the larger rectangle is 2L x 2H=4LH The area of the smaller one is LH. so does the area double or become 4 times greater?

OpenStudy (anonymous):

4 times?

OpenStudy (pfenn1):

perimeter of smaller rectangle is 2H+2L perimeter of large one is 2(2L) +2(2H) = 2(2L + 2H) so what happens to the perimeter

OpenStudy (pfenn1):

yes! area is 4 times greater when you double the lengths of the sides

OpenStudy (anonymous):

thankyouu:)!!

OpenStudy (anonymous):

so the peremeter too or is that doubled?

OpenStudy (pfenn1):

The perimeter of the larger rectangle is 2 (2H +2L) or double!

OpenStudy (anonymous):

thankss :)

OpenStudy (pfenn1):

you're welcome.